OpenStudy (anonymous):

Determine the zeros of the function x2 − 6x = -8

OpenStudy (acxbox22):

x^2-6x+8=0 (x-4)(x-2)=0 x-4=0 x-2=0 x=4 x=2 x=4,2

OpenStudy (anonymous):

thank you it help me :)

OpenStudy (acxbox22):

np